a square pyramid has a slant height of 18 centimeters and base edges of 60 centimeters find the surface area of the pyramid
1
Expert's answer
2019-05-14T10:19:17-0400

The total surface area of a pyramid is the sum of the areas of its lateral faces and its base. The base in this case is a square with side 60cm in lenght. So its area equals 60^2 = 3600 cm^2. Lateral surface area equals 1/2 * p * l, where p is a perimeter of the base and l is a slant height. So lateral surface area is 1/2 * (60+60+60+60) * 18 = 2160 cm^2. The total surface area equals 3600 + 2160 = 5760 cm^2.